Let 𝑭 = π’™π’›π’Š βˆ’ 𝒙𝒋 + π’š πŸπ’Œ. Evaluate ∭ 𝑭𝒅𝑽 𝑽 where V is the region bounded by the surfaces 𝒙 = 𝟎, π’š = 𝟎, π’š = 𝟏, 𝒛 = 𝒙 𝟐 , 𝒛 = 𝟏


1
Expert's answer
2021-06-04T11:24:20-0400
"\int\int\int_V\vec Fdv=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\displaystyle\int_{x^2}^{1}(xz\vec i-x\vec j+y^2\vec k)dzdydx"

"=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\big[\dfrac{xz^2}{2}\vec i-xz\vec j+y^2z\vec k\big]\begin{matrix}
 1 \\
 x^2
\end{matrix}dydx"

"=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\big[\dfrac{1}{2}(x-x^5)\vec i-(x-x^3)\vec j+y^2(1-x^2)\vec k\big]\begin{matrix}
 1 \\
 x^2
\end{matrix}dydx"

"=\displaystyle\int_{0}^{1}\displaystyle\int_{0}^{1}\bigg(\dfrac{1}{2}(x-x^5)\vec i-(x-x^3)\vec j+y^2(1-x^2)\vec k\bigg)dydx"

"=\displaystyle\int_{0}^{1}\bigg[\big(\dfrac{1}{2}(x-x^5)y\vec i-(x-x^3)y\vec j+\dfrac{y^3}{3}(1-x^2)\vec k\big)\bigg]\begin{matrix}
 1 \\
0
\end{matrix}dx"

"=\displaystyle\int_{0}^{1}\bigg(\big(\dfrac{1}{2}(x-x^5)\vec i-(x-x^3)\vec j+\dfrac{1}{3}(1-x^2)\vec k\big)\bigg)dx"

"=[(\dfrac{1}{4}x^2-\dfrac{1}{12}x^6)\vec i-(\dfrac{1}{2}x^2-\dfrac{1}{4}x^4)\vec j+(\dfrac{1}{3}x-\dfrac{1}{9}x^3))\vec k]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=\dfrac{1}{6}\vec i-\dfrac{1}{4}\vec j+\dfrac{2}{9}\vec k"